OpenStudy (moongazer):

If tan (theta) =5/4 and cos (theta) < 0 . Find the other trig. functions. Could someone check my answers?

OpenStudy (moongazer):

am I right that theta is in quadrant 3 ?

OpenStudy (anonymous):

Yup, so Sin (theta) < 0 as well

OpenStudy (moongazer):

Could you show your answers so that I can check my answers?

OpenStudy (anonymous):

No.. you should give your answers so that we can check.. Go ahead..

OpenStudy (moongazer):

ok

OpenStudy (moongazer):

\[\sin \theta =(-5\sqrt{41})/41\] \[\cos \theta = (-4\sqrt{41})/41\] \[\tan \theta = 5/4\] \[\csc \theta = -\sqrt{41}/5\] \[\sec \theta =-\sqrt{41}/4\] \[\cot \theta = 4/5\]

OpenStudy (moongazer):

Am I correct?

OpenStudy (anonymous):

 Just remember "All Science Teachers Crazy" so in the first quad all trig functions are positive, in the 2nd all Sine functions are positive, in the 3rd all Tangent functions are positive so in the 4th all Cosine functions are positive.